OpenStudy (anonymous):

Is the domain of y = 1/(x – 10) all real numbers.

OpenStudy (mrnood):

what happens if x=10?

OpenStudy (mrnood):

have another look at the equation if x = 10 what happens to y?

OpenStudy (anonymous):

y=0 right

OpenStudy (mrnood):

no write the equation in full, but put 10 where there is x

OpenStudy (anonymous):

ya thats what i did

OpenStudy (mrnood):

no you didn't write it out, on paper at home - or type it here - but take care. The answer is not y=0 \[y=\frac{ 1 }{ (x-10) }\] x=10 this is a simpler version of the other question you are viewing - just look carefully

OpenStudy (anonymous):

dude i keep getting zero

OpenStudy (mrnood):

(x-10) x=10 x-10 = 0 1/(x-10) = 1/0 what is 1/0?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

right???

OpenStudy (mrnood):

OK this is wher the fundamental concept lies - and why you are not getting the answer to these questions. 1/0 is not 0 it is undefined - you cannot divide by 0 - it results in an infinite number - and is therefore undefined. Therefore in any equation which has a fraction in it (such as the one in this question) the value of y is not defined when the denominator is zero In this case the denominator is zero when x = 10. But 10 is a real number, and results in an undefined value for y therefore you should be able to answer "is y defined for all real numbers"? That is the answer to this question.

OpenStudy (anonymous):

THANK YOU!

OpenStudy (phi):

Just to repeat, we do *not* allow "divide by zero". So we exclude any numbers that cause that catastrophe. In this case, we do not allow x to be 10. we say the domain (allowed x's) is all real numbers *except* x=10